ellen wants to make curtains for her kitchen.each panel takes 67and 3/4 inches of material.how many inches of material will she need if shs has three windows and each needs two panels? 3 windows with 2 panels=6 panels
Answer by tutorcecilia(2152)

3 windows with 2 panels=6 panels
6 panels X 67 and 3/4 = 6(67 + 3/4)= 402 + 4.5 = 406.5 inches